Roger makes 70 gallons of pink paint by mixing 21 gallons of red paint with 49 gallons of white paint
What part of every gallon is from red paint ?

Answer by stanbon(75887)

Roger makes 70 gallons of pink paint by mixing 21 gallons of red paint with 49 gallons of white paint
What part of every gallon is from red paint ?
--------------
Ans: 21/(21+49) = 21/70 = 3/10 = 30% red
-------------
Cheers,
Stan H.
------------